OpenStudy (jennychan12):

Find an antiderivative F of f(x) = x^2*sin(x^2) such that F(1) = 0.

OpenStudy (jennychan12):

I don't think you can use u-sub here. cuz then it'd be f(u) = u*sinu..

OpenStudy (jennychan12):

and i haven't learned trig sub or integration by parts yet.

OpenStudy (anonymous):

jenny you are a liar i think

OpenStudy (jennychan12):

-_- i haven't. but i can do integration by parts. but my teacher never taught it.

OpenStudy (anonymous):

becos that problem required fresnal integral

OpenStudy (jennychan12):

what?

OpenStudy (anonymous):

Use \(u = x^2\). Then \(x = \sqrt{u}\)

OpenStudy (anonymous):

u can use u sub on that

OpenStudy (jennychan12):

\[f(x) = x^2\sin(x^2) \]

OpenStudy (jennychan12):

if u = x^2 then du = 2xdx

OpenStudy (anonymous):

well what method are you allowed to use if you can't use trig or parts

OpenStudy (jennychan12):

that's what i don't understand. u-sub doesn't really work...

OpenStudy (anonymous):

so ur on u-sub section of the book?

OpenStudy (jennychan12):

no it's a review packet. i learned area, u-sub, trapezoid rule, and simpson's rule

OpenStudy (anonymous):

thats the way it is

OpenStudy (anonymous):

not an exponent

OpenStudy (jennychan12):

no it's \[f(x) = x^2\sin(x^2)\]

OpenStudy (anonymous):

This isn't a candidate for u sub.

OpenStudy (anonymous):

u = x^2 du = 2x dx cant use u du sub

OpenStudy (anonymous):

no place to apply a substitution

OpenStudy (anonymous):

\[ \frac{1}{2}\int \sqrt{u}\sin(u)du \]Isn't helpful.

OpenStudy (anonymous):